OpenStudy (anonymous):

I need help with coming up with my own equations for intersecting lines, parallel lines, and the same line.

OpenStudy (irishboy123):

straight lines always follows the same format in Cartesian space - ie the 2 dimensional co-ordinate system, where points in the space are identified as (a,b). a typical equation: y = mx + x m is the gradient which the ratio of how much a line rises in the vertical y axis/direction to how much it travels across the x axis/direction. this means that any of these following lines will never meet because they are parallel: y = 2x-3 y = 2x+5 y= 2x+100 they each have a gradient of +2 and will therefore never ever meet. [a line with a negative gradient falls as it travels out on the x axis. eg y = -2 +3] that's the m in y = mx + c. the c is the intercept, ie where the live touches the y axis at x = 0. so in the first example of y = 2x - 3, when x = 0, y = 3. In the second, when x = 0, y = 5. we now know that these line sit one above the other. they have the same gradient and will never ever meet because they are parallel, and one sits 2 units above the other on the y axis.